OpenStudy (anonymous):

if f(x) = 1/x what is f''(x)?

OpenStudy (anonymous):

f'(x) = -1/x^2 f''(x) = 2/x^3

OpenStudy (anonymous):

what he said

OpenStudy (anonymous):

she changed the original function

OpenStudy (anonymous):

if she asked for the 3rd derivative it would be -6/x^4

OpenStudy (anonymous):

wait are you trying to find the 40th derivative of 1/x?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

never mind lol